The number of defects in a sample of 20 parts is recorded for quality control purposes; over the last year the number of defects and their occurrence rate has been: 0 defects, 82%; 1 defect, 13%, 2 defects, 4%, 3 defects, 1%. Find the expected number of defects in the next sample of 20 parts.
Found 2 solutions by edjones, stanbon:
Answer by edjones(7569)

1*.13+2*.04+3*.01=.24
20*.24=4.8 the expected number of defects in the next sample of 20 parts.
.
Ed

Answer by stanbon(57282)

The number of defects in a sample of 20 parts is recorded for quality control purposes; over the last year the number of defects and their occurrence rate has been:
0 defects, 82%;
1 defect, 13%,
2 defects, 4%,
3 defects, 1%.
Find the expected number of defects in the next sample of 20 parts.
------------------------------------------
E(x) = 0*0.82 + 1*0.13 + 2*0.04 + 3*0.01
------------------------------------------
E(x) = 0 + 0.13 + 0.08 + 0.03 = 0.24 is the expected # of
defects in a sample of 20 parts.
============================
Cheers,
Stan H.